1000000^3x=10, 3x is the exponent.
Answer by stanbon(75887)

1000000^3x=10, 3x is the exponent.
-----
(3x)log(1000000) = log10
------
3x = 1/6
x = 1/18
-----------------
Cheers,
Stan H.